OpenStudy (anonymous):

Audrey is trying to find the equation of a line parallel to y = 2 over 3x −5 in slope-intercept form that passes through the point (−6, −1). Which of the following equations will she use? y − (−6) = 2 over 3(x − (−1)) y − (−1) = 2 over 3(x − (−6)) y − (−6) = 3 over 2(x − (−1)) y − (−1) = 3 over 2(x − (−6))

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

y = 2 over 3x −5 what is the slope of this line ? what do you know about parallel lines ?

OpenStudy (anonymous):

i know that parallel lines are two lines that lie within the same plane and never intersect and im not sure the slope would be

Nnesha (nnesha):

y = mx+ b is a slope intercept form where m is slope and b is y-intercept so what is m in this equation ??\[\huge\rm y =\frac{ 2}{ 3}x −5\]

OpenStudy (anonymous):

m= 2/3 and b= -5 right (:

Nnesha (nnesha):

yes right so parallel lines should have same slope so equation of new line should be the same :-)

Nnesha (nnesha):

and then point slope formula \[\huge\rm y-y_1 = m(x-x_1)\] where m is slope plug in (x_1 ,y_1) by order pair

OpenStudy (anonymous):

ok so 2/3 would be y1 i think and then -8 would be M is that right ?

OpenStudy (anonymous):

@Nnesha are you there?

Nnesha (nnesha):

2/3 is slope

Nnesha (nnesha):

how did you get -8 given order pair is (-6,-1) first number is x_1 and 2nd one is y_1

OpenStudy (anonymous):

oh ok sorry nnesha would it be ok if you explain it alittle better sorry im still learning about this stuff

Nnesha (nnesha):

alright so your answer choices are in point slope form which is \[\huge\rm y-y_1=m(x-x_1)\] where m is slope and \[(x_1,y_1)\] is the given order pair so just substitute values into the formula m is slope like i said parallel lines should have the same slope :-)

Nnesha (nnesha):

if slope of first line is 2/3 slope of 2nd line is also 2/3

OpenStudy (anonymous):

ohhhh ok i understand now so should the equation look like this then